The product of two consecutive whole numbers is equal to the sum of the square of the smaller number and 13.
Answer by ikleyn(52776)

.
Good style requires to put a question into the post.
~~~~~~~~~~~~~~~~~~~~~

n*(n+1) = n^2 + 13,  or

n^2 + n = n^2 + 13,


n = 13.


Check.